Determine the limit when x goes to 0 of (sin(1/x)) if it exists.


1
Expert's answer
2021-02-02T05:00:44-0500

The limit does not exist.

"h=\frac 1x, ~x\rightarrow0,~h\rightarrow \infin,"

"\lim\limits_{x\rightarrow0}sin\frac 1x=\lim\limits_{h\rightarrow\infin}sin(h)."

As "h" gets bigger, "sin(h)" keeps fluctuating between -1

 and 1. It never tends towards anything, or stops fluctuating at any point.